OpenStudy (anonymous):

Please anyone can help with my assignment please. I don’t know how to answer this our professor didn’t explain this well PROBLEMS ARE: Give the slope and intercept of each line represented by the reduce equation to the slope- intercept form 1. X-4y=8 2. 7x+y=11 3. 4x-3y=12 4. 3x+7y-6=0 Write the equation of the line determined by the slope m and y- intercept b 1. M=-4, b=7 2. M=-2/3, b=2 3. M=0, b=11 Find the equation of the line passing through the points A and B in each of the following exercises 1. A (5,1), B(1,4) 2. A(-2,3) B(7,4)

OpenStudy (anonymous):

You round, chai?

OpenStudy (anonymous):

x-4y=8y =-4y=-x+8 =y=1/4x-2 m=1/4 , and y intercept is -2

OpenStudy (anonymous):

the last set of problems requires point slope formula y2-y1=m(x2-x1)

OpenStudy (radar):

1. x-4y=8 -4y=-x+8 y=(1/4)x-2 m=1/4, y intercept at x=-2 1.m=-4, b=7 y=-4x+7 1. A(5,1), B(1,4) \[m=\Delta y/\Delta x\] m=(1-4)/(5-1))=-3/4 (y-1)=(-3/4)(x-5) y-1=-(3/4)x+15/4 y=(-3/4)x+19/4 The first one of each example, just follow the same procedure for the remaining ones

OpenStudy (anonymous):

I think the second set is just about inputting the data in y=mx+b

OpenStudy (anonymous):

yes im arround hokisan, thankyou for helping me